solve the following equation:
1/7(x-20)=1/5(x-28)

Answer by algebrahouse.com(1659)

solve the following equation:
1/7(x - 20) = 1/5(x - 28)

You can get rid of the fractions by multiplying the entire
equation by the least common denominator of the fractions, which is 35

35[1/7(x - 20) = 1/5(x - 28] {multiplied entire equation by LCD, 35}
35[1/7(x - 20)] = 35[1/5(x - 28)] {multiplied each term by 35}
5(x - 20) = 7(x - 28) {cancelled and multiplied what's left}
5x - 100 = 7x - 196 {used distributive property}
5x = 7x - 96 {added 100 to both sides}
-2x = -96 {subtracted 7x from both sides}
x = 48 {divided both sides by -2}
www.algebrahouse.com